p(x)=x^4+5x^3+4x^2-3x^2+9=(x-k)^2(ax^2+bx+c) where k is an integer and a,b,c are constants.Find the value of 5a-7b+8c

Grade:10

1 Answers

Sher Mohammad IIT Delhi
9 years ago
for x=0 9=(k^2)c , if c=1 , k=3
for x=1 16=4(a+b+1) a+b =3
for x=2 53=4a+2b+1 4a+2b=52

2a=46, a=23 , b=-20
5*46-7*(-20)+8=378
sher mohammad
b.tech, iit delhi